### v1

# Torus Knots in Adjoint Representation

Abstract
We derive a closed-form expression for the adjoint polynomials of torus knots and investigate their special properties. The results are presented in the very explicit double sum form and provide a deeper insight into the structure of adjoint invariants essential for the Vogel’s universality of Chern-Simons theory.
FIAN/TD-23/25
ITEP/TH-39/25
IITP/TH-32/25
a Lebedev Physics Institute, Moscow 119991, Russia
b
NRC “Kurchatov Institute”, 123182, Moscow, Russia
c
Institute for Information Transmission Problems, Moscow 127994, Russia
dCenter for Quantum and Topological Systems (CQTS), NYUAD Research Institute,
New York University Abu Dhabi, PO Box 129188, Abu Dhabi, UAE
Introduction.
Evaluating the HOMFLY-PT invariant of various knots and links is an important goal because of a series of applications. First of all, in accordance with [1], this invariant colored with a Young diagram $R$ is nothing but the Wilson loop average in representation $R$ in the $SU(N)$ Chern-Simons theory. Hence, studying it allows one to study pure topological effects in gauge theories. In particular, the study of adjoint representation gives information about the pure gluodynamics.
The adjoint invariants are also applied to study of the Vogel’s universality of Chern-Simons theory [2] (see also [3] for a latest review), which gives a unified description of various quantities in the adjoint sector of representation theory for all simple Lie algebras in terms of three parameters $\alpha,\beta,\gamma$, which are homogeneous coordinates of the Vogel’s projective plane. These quantities vary from the Chern-Simons partition function [4] to the Racah matrix [5] involving the adjoint representation and its descendants [6]. In particular, in [7, 8] we have constructed universal invariants for torus knots $T[2,2n+1]$, $T[3,3n\pm 1]$, and $T[4,2n+1]$. However, considering more general knots has not been possible so far.
There is also a pure knot theory application: the knowledge of the adjoint invariant allows one to evaluate the fundamental HOMFLY-PT of the reversed 2-cable knot $K$, which has a peculiar “panhandle” shape, in accordance with [9, Eq.(2)]
$H_{[1]}^{C_{2}(K)}(q,A)=\frac{1}{D_{[1]}(q,A)}\big(1+D_{\mathrm{Adj}}(q,A)H_{\mathrm{Adj}}^{(K)}(q,A)\big),$ (1)
and similarly for link [9, Eq.(33)], where $H_{R}$ denotes the (reduced) HOMFLY-PT invariant colored with the Young diagram $R$ and normalized to unknot, and $D_{[1]}$ ($D_{\mathrm{Adj}}$) is the quantum dimension of the fundamental (adjoint) representation. An important point is that the invariants are here taken in the vertical framework [9, 10].
The simplest class of knots, which we are going to study here is the torus knots $T[m,n]$. They are given by a pair of coprime numbers $m$ and $n$ (the common non-trivial divisor $l$ of $m$ and $n$ would lead to an $l$-component link). More concretely, we calculate the HOMFLY-PT invariant for these knots in the adjoint representation and in the topological framing so that formula (1) looks as
$H_{[1]}^{C_{2}(T[m,n])}(q,A)=\frac{1}{D_{[1]}(q,A)}\big(1+A^{-2(m-1)n}D_{\mathrm{Adj}}(q,A)H_{\mathrm{Adj}}^{(T[m,n])}(q,A)\big),$ (2)
where we manifestly take into account the factor of transition to the vertical framing [9]. Using the Rosso-Jones formula [11, 12] allows us to present an answer for the adjoint HOMFLY-PT polynomial of arbitrary torus knots.
Notation.
Throughout the paper, we use the notion
$\{x\}:=x-{1\over x},\ \ \ \ \ \ \ \ \ [n]_{q}:={\{q^{n}\}\over\{q\}}.$
Rosso-Jones formula.
We start with the Rosso-Jones formula [11, 12] for the (normalized) invariant of the torus knot $T[m,n]$ ($m$ and $n$ are coprime) colored by a finite-dimensional irreducible representation $R$ of the simple Lie group associated with root system $\Delta$ ($\Delta_{+}$ denotes the positive roots):
$P^{[m,n]}_{R}(q,\Delta)\ =\ \frac{q^{mn\varkappa_{R}}}{\hbox{qD}_{\!{}_{R}}(q)}\sum_{Q\in R^{\otimes m}}c_{{}_{RQ}}^{(m)}\cdot q^{-\frac{n}{m}\varkappa_{{}_{Q}}}\cdot\hbox{qD}_{{}_{Q}}(q,\Delta).$ (3)
Here
$\varkappa_{Q}=(\Lambda_{Q},\Lambda_{Q}+2\rho)$
is the eigenvalue of the second Casimir operator in the representation $Q$ given by the highest weight vector $\Lambda$,
$\hbox{qD}_{Q}(q,\Delta)=\prod_{\alpha\in\Delta_{+}}{[\left(\Lambda_{Q}+\rho,\alpha\right)]_{q}\over[\left(\rho,\alpha\right)]_{q}}$
is the quantum dimension, and the coefficients $c_{{}_{RQ}}^{(m)}$ are defined by the Adams operation: $m$-plethysm
$\hat{A}d_{m}\chi_{{}_{R}}(p_{k})\equiv\chi_{{}_{R}}(p_{mk})=\sum_{Q\in R^{\otimes m}}c_{{}_{RQ}}^{(m)}\chi_{{}_{Q}}(p_{k}),$ (4)
where $\chi_{{}_{R}}(p_{mk})$ is the character of representation $Q$. In these formulas, $\rho={1\over 2}\sum_{\alpha\in\Delta_{+}}\alpha$ is the Weyl vector. Note that expression (3) is symmetric in $m$ and $n$ (which is absolutely non-trivial). The knot invariant (3) is in the topological framing.
One of the technically direct ways to evaluate the Adams coefficients is to use the explicit formula
$c_{{}_{RQ}}^{(m)}=\sum{\psi_{R}(\Delta)\psi_{Q}(m\Delta)\over z_{\Delta}},$ (5)
where $\psi_{R}(\Delta)$ is the character of the representation $R$ of the permutation group on the conjugacy class given by the Young diagram $\Delta$, $z_{\Delta}$ is the order of automorphism of the Young diagram $\Delta$: if $\Delta=[\ldots,3^{r_{3}},2^{r_{2}},1^{r_{1}}]$, where some $r_{k}$ may be equal to zero, then $z_{\Delta}:=\prod_{k}k^{r_{k}}r_{k}!$. The Young diagram $m\Delta$ is understood as the Young diagram with all lengths of lines multiplied by $m$.
Below we obtain the answer for the adjoint HOMFLY-PT invariant (i.e. for the $A_{N-1}$ series) for an arbitrary torus knot.
$A_{N-1}$ series.
In the $A_{N-1}$ series case, all the ingredients of the Rosso-Jones formula look as follows.
The finite-dimensional irreducible representation is labeled by the Young diagram $R$, and the corresponding character is given by the Schur function $S_{R}$ [13], which is realized either as a graded polynomial of ${\rm tr}\,g^{k}$, where $g$ is the group element of $SL(N)$, or as a graded symmetric polynomial of the eigenvalues of $g$.
The Casimir eigenvalue in this case is
$\varkappa_{Q}=2\sum_{\Box_{i,j}\in Q}(j-i)-{|Q|^{2}\over N}+|Q|N.$ (6)
The quantum dimension is
$\hbox{qD}(q,A)=\prod_{i,j\in R}{\{Aq^{j-i}\}\over\{q^{h(i,j)}\}}~~,$ (7)
where $h(i,j)$ is the hook length, and the product runs over boxes of the Young diagram. It can be obtained by evaluating the Schur polynomial at the point $g=q^{2\rho}$ so that $\mathrm{tr}g^{k}={\{A^{k}\}\over\{q^{k}\}}$.
We will also need the notion of composite representation [14], which is the most general finite-dimensional irreducible highest weight representation of $SL(N)$, which is associated with the Young diagram of the form
$(R,P)=\Big[R_{1}+P_{1},\ldots,R_{l_{R}}+P_{1},\underbrace{P_{1},\ldots,P_{1}}_{N-l_{\!{}_{R}}-l_{\!{}_{P}}},P_{1}-P_{{}_{l_{\!{}_{P}}}},P_{1}-P_{{l_{\!{}_{P}}-1}},\ldots,P_{1}-P_{2}\Big],$
where $l_{P}$ denotes the number of lines in the Young diagram P. This $(R,P)$ is the first (“maximal”) representation contributing to the product $R\otimes\bar{P}$, where $\bar{P}$ denotes the conjugate representation. It can be manifestly obtained from the tensor products (i.e., as a projector from $R\otimes\bar{P}$) by formula [14]
$(R,P)=\sum_{Y,Y_{1},Y_{2}}(-1)^{l_{\!{}_{Y}}}N^{R}_{YY_{1}}N^{P}_{Y^{T}Y_{2}}\ Y_{1}\otimes\overline{Y_{2}}\,$, (8)
where the superscript “T” denotes transposition.
The adjoint representation is the simplest example of the composite representation, it is given by the Young diagram $21^{N-2}]$, which depends on $N$. We introduce the variable $A:=q^{N}$. This variable $A$ should not be confused with the second variable $A$ of the HOMFLY-PT polynomials: the HOMFLY-PT polynomial is evaluated for one and the same Young diagram for various $A_{N-1}$, while the adjoint (or, more generally, composite) representation depends on $N$ itself. Formally, we define $H_{\mathrm{Adj}}(q,A)$ via the specialization
$H_{\mathrm{Adj}}(q,A)\;\big|_{\,A=q^{N}}\;=\;J_{[2,1^{N-2}]}^{SL(N)}(q),$ (9)
for all sufficiently large $N$ [15][20]. Here $J_{R}^{SL(N)}(q)$ is the $SL(N)$ invariant. This polynomial is called uniform HOMFLY-PT polynomial in [7], and the same construction applies to any composite representation.
Our goal now is to construct the uniform (adjoint) HOMFLY-PT polynomial for the torus knots $T[m,n]$, i.e. $m$ and $n$ are coprime. Note that, in this case, constructing torus knot invariants is generally simpler than links: for instance, $\mathrm{Adj}^{\otimes 4}$ decomposes into 40 irreps, while the Adams operation generates only 17 irreps [8]. This is because the plethystic expansion of the adjoint Schur function generating the Adams coefficients is relatively simple. For instance, the 5-plethystic expansion is
$\hat{A}d_{5}S_{\mathrm{Adj}}$ $=$ $4+S_{[2^{5}1^{N-10}]}-S_{[3^{5}2^{N-9}1^{3}]}+S_{[4^{5}3^{N-8}2^{2}]}-S_{[5^{5}4^{N-7}3]}+S_{[6^{5}5^{N-6}]}$- (10)
- $S_{[32^{3}1^{N-9}]}+S_{[43^{3}2^{N-8}1^{3}]}-S_{[54^{3}3^{N-7}2^{2}]}+S_{[65^{3}4^{N-6}3]}-S_{[76^{3}5^{N-5}]}+$
$+$ $S_{[42^{2}3^{N-8}]}-S_{[53^{2}2^{N-7}1^{3}]}+S_{[64^{2}3^{N-6}2^{2}]}-S_{[75^{2}4^{N-5}3]}+S_{[86^{2}5^{N-4}]}$-
- $S_{[521^{N-7}]}+S_{[632^{N-6}1^{3}]}-S_{[743^{N-5}22]}+S_{[854^{N-4}3]}-S_{[965^{N-3}]}+$
$+$ $S_{[61^{N-6}]}-S_{[72^{N-5}1^{3}]}+S_{[83^{N-4}22]}-S_{[94^{N-3}3]}+S_{[10,5^{N-2}]}~,$
which has to be compared with just antisymmetric part of $\mathrm{Adj}^{\otimes 5}$ in [16].
The adjoint quantum dimension and Casimir exponential are
$\begin{array}[]{rclrcl}\hbox{qD}_{\mathrm{Adj}}(q,A)=\hbox{qD}_{[21^{N-2}]}(q,A)&=&\displaystyle{\{Aq\}\{A/q\}\over\{q\}^{2}},&q^{\varkappa_{\mathrm{Adj}}}&=&A^{2}.\cr\end{array}$ (12)
Uniform polynomials for torus knots $T[m,n]$.
Because of simplicity of the plethystic expansion, it is technically available to construct the knot invariant in the adjoint representation for the torus knots $T[m,n]$ with arbitrary coptime $m$ and $n$. The Adams operation generates $m^{2}+1$ irreps in this case. The general answer is quite transparent:
$\hat{A}d_{m}S_{\mathrm{Adj}}$ $=$ $m-1+$ (13)
$+$ $\underbrace{S_{[2^{m}1^{N-2m}]}-S_{[3^{m}2^{N-2m+1}1^{m-2}]}+S_{[4^{m}3^{N-2m+2}2^{m-3}]}+...+(-1)^{m+1}S_{[m+1,m^{N-m-1}]}}_{m\ terms}$-
- $\underbrace{S_{[32^{m-2}1^{N-2m+1}]}+\ldots}_{m\ terms}+$
$+$ $\underbrace{(-1)^{m}S_{[42^{m-3}1^{N-2m+2}]}+\ldots}_{m\ terms}+$
$\ldots$
$+$ $\underbrace{(-1)^{m+1}S_{[m+1,1^{N-m-1}]}+(-1)^{m}S_{[m+2,2^{N-m}1^{m-2}]}+\ldots+S_{[2m,m^{N-2}]}}_{m\ terms},$
with totally $m^{2}+1$ terms in this formula.
With these formulas for the Adams operation, constructing the adjoint knot invariants is immediate. To this end, we rewrite these formulas in the composite representation notation and notice that the Adams operation gives rise to only hook Young diagrams:
$\hat{A}d_{m}S_{\mathrm{Adj}}=m-1+\sum_{a,b=1}^{m}(-1)^{a+b}S_{([a,1^{m-a}],[b,1^{m-b}])}.$ (14)
In order to justify this formula, one can note that the coefficients $c^{(m)}_{RQ}$ in the Adams operation can be obtained by splitting the decomposition of $R^{\otimes m}$ into a sum of terms with fixed symmetric patterns given by the Young diagrams $P$ of size $m$:
$R^{\otimes m}=\sum_{P\vdash m}\pi_{P}\left(R^{\otimes m}\right).$ (15)
Then, the Adams operation gives [11][12]
$\hat{A}d_{m}(R)=\sum_{P\vdash m}\psi_{P}([m])\pi_{P}\left(R^{\otimes m}\right),$ (16)
where $\psi_{P}([m])$ is the value of the character of the permutation group ${\cal S}_{m}$ in the representation $P$ on the cycle of the maximal length (i.e. on the cyclic permutation $(1,\ldots,m)$):
$\psi_{P}([m])=\left\{\begin{array}[]{cc}(-1)^{d}&\hbox{if }P=[m-d,1^{d}]\\ 0&\hbox{otherwise}.\end{array}\right$. (19)
Now we evaluate the second Casimir eigenvalues of the representations entering (14). For the composite representation $(R,P)$ with $|R|=|P|=m$, it is generally given by the formula (see, e.g., [9]) [^3] In the generic case, the formula has the form $\kappa_{(R,P)}=2N(|R|+|P|)-{(|R|-|P|)^{2}\over N}+\kappa_{R}-\kappa_{P^{T}}.$
$\kappa_{(R,P)}$ $=$ $2Nm+\kappa_{R}-\kappa_{P^{T}}\,$,
using
$\kappa_{[a,1^{b}]}=a(a-1)-b(b+1),$
one obtains
$q^{\kappa_{([a,1^{m-a}],[b,1^{m-b}])}}$ $=$ $A^{2m}q^{2m(a+b-m-1)}.$ (20)
Similarly, the quantum dimension of the composite representation is given by the formula [17], which can be rewritten in the form
$\hbox{qD}_{(R,P)}(q,A)=\hbox{qD}_{{}_{R}}(q,Aq^{-l_{\!{}_{P}}})\,\hbox{qD}_{{}_{P}}(q,Aq^{-l_{\!{}_{R}}})\,\prod_{i=1}^{l_{\!{}_{R}}}\prod_{j=1}^{l_{\!{}_{P}}}{[N+R_{i}+P_{j}+1-i-j]_{q}\over[N+1-i-j]_{q}},$
so that
$\hbox{qD}_{([a,1^{m-a}],[b,1^{m-b}])}(q,A)={\{Aq^{a+b-1}\}\{Aq^{a-m-1}\}\{Aq^{b-m-1}\}\over\{q^{m}\}^{2}\{Aq^{-1}\}}\ \prod_{i=1}^{a-1}{\{Aq^{i+b-m-1}\}\over\{q^{i}\}}\prod_{i=1}^{b-1}{\{Aq^{i+a-m-1}\}\over\{q^{i}\}}\times$
$\times\prod_{i=1}^{m-a}{\{Aq^{-i+b-m-1}\}\{Aq^{b-i}\}\over\{q^{i}\}\{Aq^{-1-i}\}}\prod_{i=1}^{m-b}{\{Aq^{-i+a-m-1}\}\{Aq^{a-i}\}\over\{q^{i}\}\{Aq^{-1-i}\}}\prod_{i=1}^{m-a}\prod_{j=1}^{m-b}{\{Aq^{1-i-j}\}\over\{Aq^{-1-i-j}\}},$ (21)
since
$\hbox{qD}_{([a,1^{m-a}])}(q,A)={\prod_{i=0}^{a-1}\{Aq^{i}\}\prod_{i=1}^{m-a}\{Aq^{-i}\}\over\{q^{m}\}\prod_{i=1}^{a-1}\{q^{i}\}\prod_{i=1}^{m-a}\{q^{i}\}}.$
Thus, we finally come to the main result of this paper, to the explicit formula for the (reduced or normalized) uniform HOMFLY-PT polynomial of the torus knot $T[m,n]$:
$\boxed{H^{T[m,n]}_{\mathrm{Adj}}(q,A)\ =\ \frac{A^{2mn}}{\hbox{qD}_{\mathrm{Adj}}(q)}\left(m-1+\sum_{a,b=1}^{m}(-1)^{a+b}A^{-2n}q^{-2n(a+b-m-1)}\cdot\hbox{qD}_{([a,1^{m-a}],[b,1^{m-b}])}(q,A)\right),}$ (22)
with the quantum dimensions given by formulas (12) and (Uniform polynomials for torus knots $T[m,n]$.).
Inserting this answer (22) into formula (2), we immediately obtain the fundamental normalized HOMFLY-PT polynomial of the reverse 2-cable $T[m,n]$ knot [9].
Properties of uniform polynomials.
With these formulas, one can check for generic torus knot $T[m,n]$ that
  • As is usual with the Rosso-Jones formula, it looks non-symmetric in $n$ and $m$ for the torus knot $T[n,m]$, but, in fact, it is symmetric. In particular, this means that one can choose the smaller of two numbers $m$ and $n$ as an upper limit of summation in (22) to have a simpler sum.
  • Since $T[m,1]$ is unknot,
    $H^{T[m,1]}_{\mathrm{Adj}}=1.$ (23)
  • In the case of $T[m,0]$, which is the pure plethysm,
    $H^{T[m,0]}_{\mathrm{Adj}}={\hbox{qD}_{\mathrm{Adj}}(q^{m},A^{m})\over\hbox{qD}_{\mathrm{Adj}}(q,A)}.$ (24)
  • One checks that the stable uniform HOMFLY-PT homology of torus knots (in the spirit of [19]) is
    $\lim_{n\to\infty}\left({q\over A}\right)^{2n(m-1)}{\cal H}^{T[m,n]}_{\mathrm{Adj}}=\hbox{qD}_{([m],[m])}(q,A)=\hbox{qD}_{m\mathrm{Adj}}(q,A),$ (25)
    where ${\cal H}^{T[m,n]}_{\mathrm{Adj}}$ denotes the unreduced uniform HOMFLY-PT polynomial. This immediately follows from (22) at $|q|<1$, $|A|<1$.
  • One can also check that
    $H^{T[m,n]}_{\mathrm{Adj}}\Big|_{q=e^{\hbar}}=1+O(\hbar^{2}),$ (26)
    which just means that the framing is topological.
Particular specializations of the polynomial $H^{T[m,n]}_{\mathrm{Adj}}$ are:
  • The Alexander polynomial
    $\mathfrak{A}^{T[m,n]}_{\mathrm{Adj}}=H^{T[m,n]}_{\mathrm{Adj}}\Big|_{A=1}=1.$ (27)
  • The special polynomial $\sigma^{T[m,n]}_{\mathrm{Adj}}(A):=H^{T[m,n]}_{\mathrm{Adj}}\Big|_{q=1}$, which, for the series of torus knots $T[2,n=2k+1]$, $T[3,n=3k\pm 1]$, $T[4,n=2k+1]$, $\ldots$, is equal to (the signs in front of $\sigma_{[1]}^{T[m,n]}$ is, of course, not seen in the uniform special polynomial)
    $\sigma^{T[m,n]}_{\mathrm{Adj}}(A)$ $=$ $\left(\sigma_{[1]}^{[m,n]}\right)^{2},$
    $\sigma_{[1]}^{T[2,n=2k+1]}$ $=$ $-{1\over 2}A^{n-1}\Big((n-1)A^{2}-(n+1)\Big),$
    $\sigma_{[1]}^{T[3,n=3k\pm 1]}$ $=$ ${1\over 6}A^{2(n-1)}\Big((n-1)(n-2)A^{4}-2(n-1)(n+1)A^{2}+(n+1)(n+2)\Big),$
    $\sigma_{[1]}^{T[4,n=2k+1]}$ $=$ $-{1\over 24}A^{3(n-1)}\Big((n-1)(n-2)(n-3)A^{6}-3(n-1)(n-2)(n+1)A^{4}+$ (28)
    $+$ $3(n-1)(n+1)(n+2)A^{2}-(n+1)(n+2)(n+3)\Big),$
    $\ldots$
    This, indeed, coincides with the special polynomial of the torus knot $T[m,n]$ with arbitrary coprime $m$ and $n$ in the fundamental representation:
    $\boxed{\sigma_{[1]}^{T[m,n]}=(-1)^{m+1}{1\over m}A^{(m-1)(n-1)}\cdot\sum_{k=1}^{m}(-1)^{k+1}\binom{n-1}{m-k}\binom{n+k-1}{k-1}A^{2(m-k)}.}$ (29)
    The symmetry of this expression in $m$ and $n$ is absolutely non-evident, but it is symmetric.
  • The $A=q^{2}$ corresponds to the spin 1 representation of $SL(2)$, and, hence, the adjoint HOMFLY-PT polynomial becomes the colored Jones polynomial: $J^{T[m,n]}_{2}(q):=H^{T[m,n]}_{\mathrm{Adj}}\Big|_{A=q^{2}}$. Indeed,
    $\boxed{H^{T[m,n]}_{\mathrm{Adj}}\Big|_{A=q^{2}}={q^{2(m-1)(n-1)}\over 1-q^{6}}\left(1-q^{4m+2}-q^{4n+2}+q^{4(n+m)}+q^{2(m+1)(n+1)-2}(1-q^{2})\right),}$ (30)
    which coincides with the colored Jones polynomial $J_{2}$ of the torus knot $T[m,n]$. In this case, the symmetry between $m$ and $n$ is manifest.
Concluding remarks.
We presented an explicit formula(22) for the adjoint HOMFLY-PT polynomial for an arbitrary torus knot $T[m,n]$. A natural extension of the present analysis is to the case of torus links, which is considerably more involved. The complexity arises for two main reasons. First of all, the number of representations $Q$ that contribute to the r.h.s. of the Rosso-Jones formula in the case of knots is typically less than half that in the case of links. This is because the $m$-plethysm expansion of the adjoint representation gives rise to composite representations $(R,P)$ with $|R|=|P|=m$ only (and not all of possible representations even with this restriction emerge), while, in the case of $l$-component links, the sum in the Rosso-Jones formula runs over representations with $|R|=|P|$ equal to $mk/l$ with $k=1,\ldots,l$ (see [9]). An additional complication arises from the non-trivial multiplicities of representations appearing in the sum. Because of these reasons, calculating the adjoint HOMFLY-PT invariants for arbitrary links remains a future problem to solve.
Another essential problem that remains is to make a similar calculation for arbitrary knots, but for other groups. This would allow one to construct the universal invariant in the Vogel’s sense. Finally, it would be interesting to generalize the present construction to refined Chern–Simons theory, in the spirit of [18].
Acknowledgements.
We are grateful to Alexander Stoimenov and Hisham Sati, whose intelligent questions were one of the reasons that prompted us to conduct this research. The work of A.M. was partially funded within the state assignment of the NRC Kurchatov Institute, was partly supported by the grant of the Foundation for the Advancement of Theoretical Physics and Mathematics “BASIS” and by Armenian SCS grants 24WS-1C031. The work of V.K.S. is supported by Tamkeen UAE under the NYU Abu Dhabi Research Institute grant CG008